You wish to fence a circular garden of area 90 square meters. How much fence do you need?
Round your answer to three decimal places.

Found 2 solutions by Cromlix, FrankM:
Answer by Cromlix(4381)

Hi there,
Area of a circle = π x radius^2
90 = πr^2
r = √90/π
r = 5.3524
Circumference of a circle = π x diameter
Circumference = π(2(5.3524))
Circumference = 33.630 meters.
This is how much fence you need.
Hope this helps :-)

Answer by FrankM(1040)

pi * r^2 = 90
r^2 = 90/pi = 28.6479 and so
r = 5.3524.
The circumference is 2pi r or 10.7047 * pi = 33.630 meters of fence. Rounded from 33.629947.